OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w. A. w=p-21 B. w=p/21 C. w=2p/2l D. w=p-2l/2

OpenStudy (nurali):

P = 2l + 2w subtract both side by 2l and divide both side by 2 P-2l=2w w=p-2l/2

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

P=2l+2W P-2l=2W P-2l/(2)=W W=P-2l/2 Answers D

OpenStudy (mathstudent55):

@Nurali is correct, but you must be careful how you write it in one line. \(P = 2l + 2w\) \(P - 2l = 2w\) \(\dfrac{P - 2l}{2} = w\) \(w = \dfrac{P - 2l}{2}\) \(w = (P - 2l)/2\) ( which is different from \(w = P - 2l/2\) ). Remember that a fraction means both division and grouping.